# The process of polarisation as a loss of dimensionality: measuring changes in polarisation using Singular Value Decomposition of network graphs

### v1

@@ Abstract
The increasing polarisation in our societies has been the focus of much research and is a major international concern. Current approaches to defining and detecting polarisation largely rely on finding evidence of bimodality in social networks or voter opinion surveys. Much of this research is based on the USA’s Republican and Democrat parties, which can be difficult to apply to political systems with more than two parties or blocs. This approach also makes it hard to detect temporal trends in polarization, as the results usually fall into a binary of polarised or non-polarised; it is difficult to robustly show that subsequent increases in the bimodality of a polarised distribution are meaningful changes.
Our work is aligned with Baldassarri and Gelman’s theory that polarisation should be defined as increasing correlation between positions in the ideological field, which reduces political pluralism. We also draw from post-structuralist work which argues that polarisation is the process of both the ideological fiend and the material society being segregated into two poles, as in cases of apartheid. This means that in order to measure the polarisation occurring in a society, it would be beneficial to be able to assess social networks directly.
To measure polarisation in the social network, we use Random Dot Product Graphs to embed social networks in metric spaces. In the case of a social network, the embedded dimensionality corresponds to the number of reasons that two people may form a social connection. A decrease in the optimal dimensionality for the embedding of the network graph, as measured using truncated Singular Value Decomposition of the graph adjacency matrix, is indicative of increasing polarisation in the network.
We apply this method to the communication interactions among New Zealand Twitter users discussing climate change issues, from 2017 to 2023. We find that the discussion is becoming more polarised over time, as shown by a decrease in the dimensionality of the communication network. Second, we apply this method to discussions of the COP climate change conferences, showing that our methods agree with other researchers’ detection of polarisation in this space. Finally, we use networks generated by stochastic block models to explore how an increase of the isolation between distinct communities, or the increase of the predominance of one community over the other, in the communication networks are identifiable as polarisation processes.
@@
polarization , ideological polarization , political polarization , social complexity , climate change , random dot product graphs , graph dimensionality , singular value decomposition
@@ 1 Introduction
Social and political polarisation is an issue of increasing concern in the international community, since it is associated with severe social divisions and weakening of the democratic consensus [1]. Much of the research on polarisation is directed at the USA, particularly in the wake of the Trump presidency, but concern about polarisation is also present in a wide range of countries [2] [3] [4] [5].
@@
Data-based definitions of polarisation primarily focus on the presence of a bimodal division in the distribution of ideological or social affiliations, such as the spectrum of left- to right-wing politics; bimodality is assessed by inspection (e.g. no overlap between the ideological groups)[6] or tested for significance using tests such as Hartigan’s Dip test [7][8]. However, this approach can struggle when assessing multi-party democracies that do not have a clear left- and right-wing split [9]. When investigating social networks, bimodality is expressed as strong in-group/out-group divisions [10], and is often characterised by hostility between the two groups; it is common to consider this hostility as a sign of polarsation, not just the differences in policy positions between the two poles (usually the Republican and Democrat parties in the USA) [11] [12]. On social media, these hostile divisions usually take the form of ”echo chambers” that focus on one set of political views and exclude all others [13] [14].
@@
Another common axiom in current polarisation studies is that the polarisation is maximised when the groups are of equal size, as well as strongly divided [15]. This is intended to help tell polarisation conflicts apart from other major social conflicts, such as conflicts over wealth inequality. However, a consequence of this axiom is that polarisation cannot become severe if one of the groups is a small minority (e.g. refugees); given that some major social divisions are based on opposition to small minority groups, we are interested in whether there are methods for measuring polarisation that do not require this axiom.
@@
Computational data science approaches to measuring polarisation are popular, especially for their speed when analysing millions of records from social media services. There are fast and simple algorithms for detecting communities that can then be investigated for polarisation, such as the Louvain method [16]. Network data is usually processed using a dimension reduction algorithm, such as Principal Component Analysis or Canonical Correspondence Analysis, in order to create a single dimension that can be evaluated for bimodality. Measuring latent positions empirically means projecting them in lower dimensional spaces, and then assess the resulting first dimension for bimodality [17] In other cases, clustering algorithms are more appropriate than bimodality tests; there are a number of ways of measuring distances between the clusters, which can indicate polarisation between the groups if the clusters also have low internal variation [18].
@@
There are criticisms of the focus on bimodality in polarisation research. Baldassarri and Gelman [19] argue that focus on radicalisation relies on correctly selecting the ideological dimension that is the centre of the polarisation; incorrect selection can result in failing to find polarisation when it is present. They instead propose a definition of polarisation as increasing correlation in the ideological space, in comparison to an integrated society which ”is not a society in which conflict is absent, but rather one in which conflict expresses itself through nonencompassing interests and identities”. The process of increasing correlation between ideological positions reduces political pluralism and restricts possible ideological opinions, until maximum correlation is reached and an oppositional binary is created. They did not find that there was increasing correlation in the ideological field of US-American voters pre-2004. However, subsequent research by Kozlowski and Murphy [20] using the same methods found that correlation rapidly increased between 2004 and 2016. They noted that the increase in polarisation was strongest in the domains of economics and civil rights issues, rather than in the domain of moral issues that the ”culture war” framing of polarisation may suggest. Similarly, DellaPosta [21] conceptualises polarisation as similar to an oil spill, with the increasing correlation in ideological positions spreading polarisation to previously apolitical members of society. The article analyses how the ”belief network” of US-American politics has changed over time, concluding that the network has developed clusters which have reduced the prevalence of cross-cutting ideological positions; this means that pluralism has decreased and polarisation has increased.
@@
Incorporating post-structuralist political theory allows us to expand on this understanding of polarisation as correlation. Ernesto Laclau & Chantal Mouffe [22] begin from the same understanding of pluralism and polarisation as scholars such as Baldassarri and Gelman, but they develop this concept beyond just they ideological field. They argue that ”In a colonized country, the presence of the dominant power is every day made evident through a variety of contents: differences of dress, of language, of skin colour, of customs […] the colonizer is discursively constructed as the anti-colonized.” (p. 128). Two poles are constructed that are mutually exclusive and have nothing in common, sustained by segregation in all layers of society (e.g. the South African regime of racial apartheid). Notably, the polarisations that they examine do not occur primarily in the division between political parties, but along fault lines such as race and ethnicity, economic class, the urban-rural divide, and the division between coloniser and colonised.
@@
In this paper we present a novel method of measuring polarisation that follows from both Laclau & Mouffe and Baldassarri & Gelman, namely using Singular Value Decomposition to determine the correlation structure of social networks. Using social networks allows us to capture interactions between people that are not explicitly political — being neighbours, sharing a workplace, etc — but which become politicised and segregated during extreme polarisation. As such, we are able to determine whether correlation is increasing not just among possible political positions, but whether it is increasing among social determinants of interaction as well. Our method gives a value that corresponds to the network’s capacity for complexity. Incorporating these additional layers of a society into the analysis should make it easier to detect whether polarisation is occurring.
@@ 2 Methods
We represent the conversation happening on a social platform (Twitter, Facebook, Instagram, etc.) as a network. Each user that took part in the conversation is mapped to a node. We add an edge between two nodes if the two respective users have communicated in the time window of the observation. Depending on the chosen social platform considered and the specific research question, a communication can be given by a reply, a mention, a quote, a repost/retweet/share, or a set of these. These networks can be directed (as is more common) or undirected. Here, we consider communication networks only as unweighted graphs, although the generalization to weighted graphs doesn’t present fundamental challenges.
@@ 2.1 Network modelling
Having established these networks, we model them as Random Dot Product Graphs (RDPGs) [23]. RDPGs are used instead of other graph embeddings because their optimal dimensionality is established a priori to the analysis, so it is independent of the network’s size.
@@
In inference tasks, starting from an observed graph, the goal is to estimate the position of the nodes in the latent spaces, given the interaction structure of the network. We do not parametrise the network for this analysis. For a fixed dimension $d$ of the two latent spaces, this is achieved by a $d$-truncated Singular Value Decomposition as follows.
Let $A$ be the adjacency matrix of $G$. Let $A=U\Sigma V^{\prime}$ be a singular value decomposition of A, so that $U$ and $V^{\prime}$ are orthogonal matrices, and $\Sigma$ is the diagonal matrix which $i$-th entry is the $i$-th singular value of $A$ (sorted in decreasing order). Notice that in general $U$ and $V^{\prime}$ are only identifiable up to orthogonal transformations (any rotation of them would keep the dot product constant, so they would determine the same graph). Denoting $M|_{k}$ the truncation of a matrix $M$ to its first $k$ columns, for any $d$, the two matrices $\hat{L}=U|_{d}\sqrt{\Sigma}$ and $\hat{R}=\sqrt{\Sigma}\left(V|_{d}\right)^{\prime}$ determine a rank-$d$ optimal approximations of $A$. That is, $\hat{L}\hat{R}=\hat{A}$ minimizes the Frobenius distance to $A$ between all the rank-$d$ matrices.
@@
An a-priori optimal choice for $\hat{d}$ can be obtained from, $\Sigma$, the sorted sequence of singular values of the network’s adjacency matrix $A$. Various methods exist. Here we adopt the elbow method presented in [24]. The elbow method identifies the most likely change point in the sequence of values of $\Sigma$ by sequentially fitting two Gaussian distributions with independent mean, and equal variance. One Gaussian distribution is fitted to the largest $d$ singular values, and the other to the smallest $K-d$. Then, $\hat{d}$ is the value of $d$ that maximise the sum of the log-likelihoods of the two distributions.
Notice that $\hat{d}$ is robust to network size.
@@
Given a network, we can assess its graph complexity by computing its SVD entropy, which is again based on $\Sigma$. A network has higher SVD entropy when many of its singular vectors are highly important for its structure, meaning that the network can not be efficiently compressed. This commonly read as an indication of high network complexity [25], and it is related (although not in a linear nor straightforward way) to its dimension. We normalise the SVD entropy using Pielou’s evenness [26], so that the results do not depend on the network size.
In particular, let $\Sigma$ be the sequence of singular values of a network’s adjacency matrix $A$. The nuclear norm of $A$ is given by the sum total of $\Sigma$ (that is, the sum of all singular values). We define the normalized values $s_{i}=\frac{\sigma_{i}}{\|A\|_{*}}$, where $i\in\{1,....N\}$ and $\sigma_{i}$ is the $i$-th singular value.
@@
2.2 Polarization
We define a process of polarization as the loss of dimensionality of a graph observed in time. Namely, we find the optimal RDPG embedding dimension $\hat{d}$ at multiple time points. If $\hat{d}$ decreases over time then we argue that the network has become more polarised during that time. This is based on the same principles as the view that the process of polarisation is one of increasing correlation. The dimensions of $\hat{d}$ are all uncorrelated; as such reduction in $\hat{d}$ corresponds to a reduction in uncorrelation in the social network graph, which is an increase in correlation.

### v3

# The process of polarisation as a loss of dimensionality: measuring changes in polarisation using Singular Value Decomposition of Random Dot Product Graphs

@@ Abstract
In this paper we present new methods that extend Baldassarri and Gelman’s theory of polarisation. Baldassarri and Gelman show that it is useful to define polarisation as increasing correlation between positions in the ideological field, which reduces political pluralism. We also draw from post-structuralist work which argues that this correlation extends to elements of culture beyond the ideological, and deliberate development of these correlations is a feature of polarised regimes such as apartheid.
To measure polarisation in social networks, we use Random Dot Product Graphs to embed social networks in metric spaces. Singular Value Decomposition of a social network provides an embedded dimensionality which corresponds to the number of uncorrelated dimensions in the network. Each uncorrelated dimension in a social network represents a part of that society which allows two people from different groups to form a social connection, such as living in a racially integrated neighbourhood. A decrease in the optimal dimensionality for the embedding of the network graph means that the dimensions in the network are becoming more correlated, and therefore the network is becoming more polarised.
We apply this method to the communication interactions among New Zealand Twitter users discussing climate change issues, from 2017 to 2023. We find that the discussion is becoming more polarised over time, as shown by a decrease in the dimensionality of the communication network. Second, we apply this method to discussions of the COP climate change conferences, showing that our methods agree with other researchers’ detection of polarisation in this space. Finally, we use networks generated by stochastic block models to explore how an increase of the isolation between distinct communities, or the increase of the predominance of one community over the other, in the social networks decrease the embedded dimensionality and are therefore identifiable as polarisation processes.
@@
polarisation , ideological polarisation , political polarisation , social complexity , climate change , random dot product graphs , graph dimensionality , singular value decomposition
@@ 1 Introduction
Social and political polarisation is an issue of increasing concern in New Zealand [1], as well as a wide range of countries [2]. [3] [4] [5] [6]. While it was initially thought that New Zealand had avoided the populist takeover seen in countries such as the USA [7], radical changes in the government makeup in both 2020 and 2023 have since called this into question. This paper will explore existing methods of measuring polarisation and propose a new measurement based on Singular Value Decomposition of social Random Dot Product Graphs.
New Zealand was considered to have escaped the development of polarisation which had developed in similar nations, particularly with the 2016 election of Donald Trump in the USA [7]. Analysis of voter polarisation in NZ from 2009-2018 showed little evidence of polarisation in that period [8]. Topic modelling of political party manifestos showed similarities between the manifestos of parties on the left and right [9], suggesting that despite their rivalries the parties are not deeply divided. Tan [10] argues that the lack of bipartisanship between NZ’s left and right indicates that the parties are polarised; however, they found that the perceived distances between left and right parties, and between those parties and voters, were stable over time, suggesting that polarisation is not increasing. Satherley et al. [11] find that there is a high level of stability in the partisanship of NZ voters (i.e. they do not frequently change which party they support) in years up until 2017 and note that this creates a risk of polarisation if voters stay committed to a party that becomes increasingly extreme. In another paper, Satherley et al. [12] find that in analysis of polarisation between the Labour, Green, and National parties, NZ European voters and voters of high socio-economic status exhibit more polarisation in their party preferences than other groups. Stanley et al. [13] find that climate change is not a polarising issue for New Zealand, with the voting public largely accepting its existence.
This history of political stability highlights how irregular the NZ election results in 2020 and 2023 have been, with significant implications for potential polarisation. The 2020 election saw the Labour party win more than 50% of the vote, the first outright majority since the change to Mixed Member Proportional government in 1996, and saw the right-wing ACT minor party increase their vote share from 0.5% to 7.58%. The covid-19 pandemic has subsequently become a focus for disinformation producers, prompting concerns about an increase in polarisation since 2020 [14] [15] [16]. The 2023 election gave further new results; despite the return to the right-wing, both the NZ First and ACT parties were required for the National party to form a coalition, creating the first instance of the National party having to negotiate with two coalition partners of this size. The 2023 Coalition Government has rapidly removed Māori language official Government documents [17], and removed references to the Treaty of Waitangi (a constitutional document that forms the basis of the relationship between the indigenous Māori people and the British settlers) from the law [18]. The ACT party have also introduced a parliamentary bill to redefine the “Treaty Principles” in ways that remove references to Māori. The Waitangi Tribunal (the permanent commission of inquiry tasked with investigating governmental breaches of the Treaty of Waitangi) has argued that “the Crown’s [Government’s] actions ‘threaten to disfigure or rupture’ the Crown–Māori relationship and ‘could set back the foundational relationships of Aotearoa New Zealand for decades’” [19]. A rise in general racist sentiment, as well as racist hate crimes and violence, has been documented [20]. As such, we think that questions about whether the polarisation may be happening along an axis other than party affiliation are particularly relevant in the current political atmosphere, since there is a risk that these policies will create a polarisation between Māori and pākehā (descendants of British settlers).
@@
Polarisation scholarship in New Zealand is particularly influenced from research in the USA, so it is useful to cover some key analyses of their system as well as general data-based polarisation measurement methods. Heatherington [21] famously found that the Republican and Democrat parties have no overlap in their policies or voting patterns, finding a bimodal division by inspection (e.g. no overlap between the ideological groups). Most subsequent research uses significance using tests such as Hartigan’s Dip test [22] [23]. When investigating social networks, bimodality is expressed as strong in-group/out-group divisions [24], and is often characterised by hostility between the two groups; it is common to consider this hostility as a key sign of polarisation, in addition to the differences in policy positions between the two poles (usually the Republican and Democrat parties in the USA) [25] [26]. On social media, these hostile divisions usually take the form of ”echo chambers” that focus on one set of political views and exclude all others [27] [28].
We have some specific concerns about the use of bimodality in detecting polarisation. The most important is that use of hypothesis tests can be used to detect when a distribution is polarised, but after the first significant result it is difficult to robustly show that any further increase in polarisation is also significant. That is to say, while subsequent results may also be statistically significant, it is difficult to show that the increase from the first result to the second is itself significant. This means that we are not confident that studies which use bimodality are able to claim that We are therefore interested in developing measurements of polarisation that do not rely on significance testing in this way. A secondary concern is that this approach can struggle when assessing multi-party democracies that do not have a clear left- and right-wing split [29]. In the New Zealand context, the NZ First party makes it difficult to clearly reduce the field to left and right blocs, as they have formed coalition governments with both the Labour and National parties many times. As such, finding a method of measuring polarisation that can better handle this type of situation would be beneficial to researchers in New Zealand and similar countries.
@@
Another common aspect of definitions of polarisation is that the polarisation is maximised when the groups are of equal size, as well as strongly divided. This was first proposed by Esteban & Ray [30]. Taking group size into account is intended to help tell polarisation conflicts apart from other major social conflicts, such as conflicts over wealth inequality. However, a consequence of this theoretical approach is that sets a rigid upper limit on how severe a polarisation can become if one of the groups is a small minority. This concerns us because of issues such as potential polarisation against groups such as refugees; they are by definition a small proportion of the population but are still considered to be a potential cause of polarisation, especially in countries that are seeing influxes of refugees due to war or climate change. We are interested in finding definitions of polarisation, and tools for measuring it, that are not as affected by the size of the groups.
@@
Computational data science approaches to measuring polarisation are popular, especially for their speed when analysing millions of records from social media services. There are fast and simple algorithms for detecting communities that can then be investigated for polarisation, such as the Louvain method [31]. In some cases, clustering algorithms are used to measure polarisation; there are a number of ways of measuring distances between the clusters, which can indicate polarisation between the groups if the clusters also have low internal variation [32]. However, this has similar issues to bimodality testing insofar as it becomes difficult to interpret when there are more than two clusters. It is possible to argue that three or more groups can all be polarised against each other, but we are unconvinced by this; as the number of polarised groups increases, so does the number of dimensions required for their clusters to all be equidistant from each other, which increases the complexity of the space so much that it becomes more complex than what we would consider to be “polarised”.
In computational analyses focussed on bimodality instead of clustering, network data is usually processed using a dimension reduction algorithm, such as Principal Component Analysis or Canonical Correspondence Analysis, in order to create a single dimension that can be evaluated for bimodality. Measuring latent positions empirically means projecting them in lower dimensional spaces, and then assess the resulting first dimension for bimodality [33]. These analyses do not treat the correlation structure delivered by the PCA or CCA as informative in and of itself; they are simply used to discover the largest principal component or axis (i.e. the largest dimension). An advantage of this approach is that the issue driving polarisation is generated from the data, rather than researchers presuming what it is and risking choosing incorrectly. However, it does not entirely mitigate the risk of choosing the wrong dimension, since it presumes that polarisation is happening along the first dimension, and it therefore will still miss cases in which polarisation is happening along one of the other dimensions.
@@
In order to comprehensively address the issue of correct dimension selection, Baldassarri and Gelman [34] propose a definition of polarisation as increasing correlation in the ideological space, i.e. that pairs of ideological issues asked about in the American National Election Study become more correlated as a result of polarisation, reducing political pluralism and restricting possible ideological opinions, until maximum correlation is reached and an oppositional binary is created. This polarisation stands in comparison to an integrated, non-polarised society which ”is not a society in which conflict is absent, but rather one in which conflict expresses itself through nonencompassing interests and identities”. In their tests of Pearson correlations of pairs of issues in the American National Election Study, they did not find that there was increasing correlation in the ideological field of US-American voters pre-2004. However, subsequent research by Kozlowski and Murphy [35] using the same methods found that correlation between pairs of ideological issues rapidly increased between 2004 and 2016. They noted that the increase in polarisation was strongest in the domains of economics and civil rights issues, rather than in the domain of moral issues that the ”culture war” framing of polarisation may suggest. Similarly, DellaPosta [36] conceptualises polarisation as similar to an oil spill, with the increasing correlation in ideological positions spreading polarisation to previously apolitical members of society. The article analyses how the ”belief network” of US-American politics has changed over time, concluding that the network has developed clusters which have reduced the prevalence of cross-cutting ideological positions; this means that pluralism has decreased and polarisation has increased.
@@
Incorporating post-structuralist political theory allows us to expand on this understanding of polarisation as correlation. Ernesto Laclau & Chantal Mouffe [37] begin from the same understanding of pluralism and polarisation as scholars such as Baldassarri and Gelman, but they develop this concept beyond just they ideological field. They argue that ”In a colonized country, the presence of the dominant power is every day made evident through a variety of contents: differences of dress, of language, of skin colour, of customs […] the colonizer is discursively constructed as the anti-colonized.” (p. 128). This is to say that it is possible for elements of a society’s culture that are not explicitly political to become associated with ideologies and drawn into the correlation, further reducing opportunities for political pluralism. Two poles are constructed that are mutually exclusive and have nothing in common, sustained by segregation in all layers of society (e.g. the South African regime of racial apartheid). Notably, the polarisations that they examine do not occur primarily in the division between political parties, but along fault lines such as race and ethnicity, economic class, the urban-rural divide, and the division between coloniser and colonised. The starkest demonstration of this principle is in the actions of the 1970s Argentinian military dictatorship, which banned Venn diagrams from being taught in primary schools because they ”were feared to encourage subversive models of collectivity” [38].
@@
In this paper we present a novel method of measuring polarisation that follows from both Laclau & Mouffe and Baldassarri & Gelman, namely using Singular Value Decomposition to determine the correlation structure of social networks. Using social networks allows us to capture interactions between people that are not explicitly political — being neighbours, sharing a workplace, etc — but which become politicised and segregated during extreme polarisation. As such, we are able to determine whether correlation is increasing not just among possible political positions, but whether it is increasing among social determinants of interaction as well. Our method gives a value that corresponds to the network’s capacity for complexity, and is inversely related to its level of polarisation. Incorporating these additional layers of a society into the analysis should make it easier to detect whether polarisation is occurring.
@@ 2 Methods
We represent the conversation happening on a social platform (Twitter, Facebook, Instagram, etc.) as a network. Each user that took part in the conversation is mapped to a node. We add an edge between two nodes if the two respective users have communicated in the time window of the observation. Depending on the chosen social platform considered and the specific research question, a communication can be given by a reply, a mention (”tagging” them in a post), a quote, a repost/retweet/share, or a set of these. These networks can be directed (as is more common) or undirected. Here, we consider communication networks only as unweighted graphs, although the generalization to weighted graphs doesn’t present any fundamental challenges.
@@ 2.1 Network modelling
Having established these networks, we model them as Random Dot Product Graphs (RDPGs) [39]. RDPGs are used instead of other graph embeddings because their optimal embedding dimension (see below) is established a priori to the analysis, so it is independent of the network’s size.
@@
In inference tasks, starting from an observed graph, the goal is to estimate the position of the nodes in the latent spaces, given the interaction structure of the network. We do not parametrise the network for this analysis. For a fixed dimension $d$ of the two latent spaces, this is achieved by a $d$-truncated Singular Value Decomposition as follows (full description of SVD can be found in Noble [40].
Let $A$ be the adjacency matrix of $G$. Let $A=U\Sigma V^{\prime}$ be a singular value decomposition of A, so that $U$ and $V^{\prime}$ are orthogonal matrices, and $\Sigma$ is the diagonal matrix whose $i$-th entry is the $i$-th singular value of $A$ (sorted in decreasing order). Notice that in general $U$ and $V^{\prime}$ are only identifiable up to orthogonal transformations (any rotation of them would keep the dot product constant, so they would determine the same graph). Denoting $M|_{k}$ the truncation of a matrix $M$ to its first $k$ columns, for any $d$, the two matrices $\hat{L}=U|_{d}\sqrt{\Sigma}$ and $\hat{R}=\sqrt{\Sigma}\left(V|_{d}\right)^{\prime}$ determine a rank-$d$ optimal approximations of $A$. That is, $\hat{L}\hat{R}=\hat{A}$ minimizes the Frobenius distance to $A$ between all the rank-$d$ matrices. ${\Sigma}$ is truncated to the first $d$ diagonal elements; $\hat{L}$ has $N$ rows and $d$ columns, and $\hat{R}$ has $d$ rows and $N$ columns.
@@
An a-priori optimal choice for $\hat{d}$ can be obtained from, $\Sigma$, the sorted sequence of singular values of the network’s adjacency matrix $A$. Various methods exist. Here we adopt the elbow method presented in [41]. The elbow method identifies the most likely change point in the sequence of values of $\Sigma$ by sequentially fitting two Gaussian distributions with independent mean, and equal variance. One Gaussian distribution is fitted to the largest $d$ singular values, and the other to the smallest $K-d$. Then, the optimal $\hat{d}$ is the value of $d$ that maximise the sum of the log-likelihoods of the two distributions.
Notice that $\hat{d}$ is robust to network size, i.e. its value is determined by the complexity of the network rather than the size of the network [39].
@@
Given a network, we can assess its graph complexity by computing its SVD entropy, which choose this instead of other entropy measures because it is based on $\Sigma$. A network has higher SVD entropy when many of its singular vectors are highly important for its structure, meaning that the network cannot be efficiently compressed. This is commonly read as an indication of high network complexity [42], and it is related (although not in a linear nor straightforward way) to its dimension. We normalise the SVD entropy using Pielou’s evenness [43], so that the results do not depend on the network size.
In particular, let $\Sigma$ be the sequence of singular values of a network’s adjacency matrix $A$. The nuclear norm of $A$ is given by the sum total of $\Sigma$ (that is, the sum of all singular values). We define the normalized values $s_{i}=\frac{\sigma_{i}}{\|A\|_{*}}$ where $\|A\|_{*}$ is the Frobenius norm of $A$, and therefore $i\in\{1,....N\}$ and $\sigma_{i}$ is the $i$-th singular value.
@@
2.2 Polarisation
We define a process of polarisation as the loss of dimensionality of a graph observed in time. Namely, we find the optimal RDPG embedding dimension $\hat{d}$ at multiple time points. If $\hat{d}$ decreases over time then we argue that the network has become more polarised during that time. This is based on the same principles as the view that the process of polarisation is one of increasing correlation. The dimensions of $\hat{d}$ are all uncorrelated; as such reduction in $\hat{d}$ corresponds to a reduction in uncorrelation in the social network graph, which is an increase in correlation.